# Counterfactual Effect Generalization: A Combinatorial Definition

### v4

@@ 6.2 Counterfactual and Correlated Outcome Prediction
While the concepts above can give researchers larger control over the generalizability of learning solutions, we started with the goal of studying the EV of counterfactual predictions. This case is show in Fig.5(b). We take this to be the prediction of effect differences, $\Delta y_{ij}$, from covariate contrasts, $x_{i}-x_{j}$, for all validation unit pairs. The figure illustrates that everyday algorithms are clearly ineffective in this case. The intuition, and why the problem is rarely framed this way, was articulated above: using factor differences leads to samples with increased pairwise overlaps and decreasing permutations, decreasing the capability of algorithms to generalize effectively. The figure shows, however, that models can generalize effectively if these overlaps are considered carefully, as in square enumerations. This echoes the increased performance observed over correlated data, and reveals a relationship between these two problems.
@@ 6.3 Causal Effect Estimation and Black-box Explanations
Supervised solutions are highly tuned to generalizability. We now consider implications to causal techniques, which, instead, emphasize biasedness and estimation under non-i.i.d. conditions. Fig.4(d) and Fig.6(d) illustrate two distinct ways to transverse enumerated squares. Each individual square column corresponds to sets of effect observations with the same ordinal statistics and thus similar confounding level and variance, Proposition.1. Distinct square transversals reproduce EV increasing (Definition.2) and CF non-increasing (Definition.3) sequences. Imagine we place squares side-by-side, and that each square cell contains a sample unit. We transverse squares from left-to-right and top-to-bottom. The right-ward sequence of columns is, however, generated in one of two ways: across squares (a square transversal, like in the previous sections) and within squares (a square-column transversal). Using input from these two transversals should make estimators behave asymmetrically in respect to EV and CF. A square transversal corresponds to an order with increasing $EV$ and minimal $CF$ increase. A square-column transversal corresponds to an order with decreasing $CF$ and minimal $EV$ increase. The enumeration production rule for a column $d_{t}$, at time $t$, for each case is also shown in the figures, $d_{1}=1$. Fig.4(c) shows how ACC changes for supervised systems under random, square, and square-column transversals. Fig.6(c) shows how ACC changes, instead, for contemporary causal effect estimators and explaining systems (discussed below). Accuracy is, in this case, sum of squared differences between estimated and ground truth single-variable effects (for all variables). For a set of enumerated squares, for references $x_{0},x_{1},...,x_{t}$ at time $t$, accuracy is calculated for populations $x_{0},x_{1},...,x_{t}$. Lines correspond to mean ACC (and errors) of the best performing algorithm and parameter set (leader) across 1000 runs. Both transversals (square and square-column) supply systems with equally-represented populations but their performance is symmetric in respect to ACC. The random sample ordering strikes a balance between these extremes.
@@
Fig.6(a) shows performance of 3 widely used causal effect estimators and the IME explainer  [Lundberg and Lee, 2017, Lundberg et al., 2019] with random (top) and square-column (bottom) orders for three levels of sample correlation (correlated, square-column transversal)999with $n<100$, omitted datapoints are due to regression non-convergence.. The first estimator is the recent Deconfounder  [Wang and Blei, 2020](Sect. 3.1, linear Bayesian factor model fit with Variational Bayes, logistic outcomes, and Normal priors). The second is g-formula  [Chatton et al., 2020], an efficient solution popular in Epidemiology (Sect. 4.4 Non-parametric Effect Estimation). IME uses the Shapley value calculated from the output of all supervised methods in the last section. A propensity score matching estimator  [Rosenbaum and Rubin, 1983] is included for its popularity and baseline significance in counterfactual solutions. These algorithms (causal and explainer) make different assumptions. The first focuses on biasedness, the second combines (biased and often heuristic) predictive regressions for model selection. As expected, causal effect estimators perform well across correlated cases, compared to the explainer, which makes i.i.d. assumptions, Fig.6(a). Effect estimators loose accuracy with increasing correlation, however, as increasingly less data is uncorrelated in samples and used by these methods.
@@
Squares were used here as means to represent sets of effect observations for a unit $x_{0}$, their permutations in samples, and thus define non-parametric statistics for the effects of sample factors. Our initial definition of effects, Eq.(3), led to consequent definitions and interpretations for the EV, Definition.2, and CF, Definition.3, of counterfactual observations and samples. The unit of analysis here is $\Delta y$, an outcome difference observation, and not the outcomes of regressions. Permutations are not random sample intake orders, but ’observed’ permutations in samples that fulfilled a set of combinatorial and gap conditions, offering additional guarantees over effect estimates. Eq.(3) thus uses the complete set of $m$-way observed differences in a sample to permute observed counterfactual outcomes and derive effect estimates. The full set of variations (EV) ${\Pi}_{n}(Z\/{a/})$ is the full set of EV-increasing sequences in samples (or their ’external variations’) for sample populations, which is often limited by sample sizes $n$, or even be asymptotically unobservable for some treatment-background assignments.
@@
Two fundamental problems with both the previous explaining and sensitivity analysis approaches are the need for parametric forms for outcomes, and the assumption of i.i.d. variables  [Aas et al., 2021]. The latter is also a problem for standard non-parametric tests and generic permutation-based tests  [Hoeffding, 1948, Yamato and Maesono, 1986]. Fig.6(a) shows that correlations introduce considerable biases into these methods’ estimates of variable importance. Fig.6(b) shows, however, that both methods perform well under square enumerations. Performance increases in both fronts (effect estimation and importance attribution) in this case. In fact, in small samples, the latter dominate recent methods designed specifically for causality and correlated data (upper-right panels). This illustrates that, when samples are limited and effect estimates not externally valid, solutions aimed at generalization (given their inputs are not biased) can offer gains in causal effect estimation. The Shapley (teal curve) and Sobol indices (black) largely coincide under these conditions, as variance becomes an unbiased indicator of importance and EV, Eq.(4). This illustrates that simple estimators (which do not involve over 20 supervised methods, repeated many times for explanations) can also become effective under square transversals. Notice that both the Deconfounder and g-formula are highly-tuned to the generative models used, as both rely on logistic regressions. This exemplifies the attraction of non-parametric insights, and that, even when models are correctly specified, there is room for gains with the proposed framework.
@@
Altogether, these results suggest that sample size increases alone can not lead to increases in the generalizability of algorithms. The number of EV-increasing sequences in the sample play a key role. At the same time, the number of CF non-increasing play a role for causal effect estimation (or other tasks related to the estimation of individual factor effect). This observation is in direct contrast with, and suggests significant penalties for, the i.i.d. assumptions of black-box and bootstrap sampling approaches. These methods can observe gains and control relevant statistical requirements by sampling squares, instead of individual observations.
@@ 6.4 Square Gap and Column Statistics
The panels in Fig.7 (a-l) illustrate the proposed empirical error decomposition for counterfactual effect observations, Eq.(9), in enumerated sample squares. These diagrams depict the mean empirical error among pairs, $(\Delta_{ij}y-\Delta_{uv}y)^{2}$, (y-axis) in their $d$-th column (x-axis). These are observed errors among sample units’ outcomes (in the same square and column), and not effect estimates. The top row, Fig.7(a-d), shows errors in the previous cases (balanced, unbalanced, both transversals). Larger $d\leq m^{\prime}$, lead to pairs exposed to increasing unobserved confounders, and variance in effects, for a fixed factor and square reference unit $x_{0}$. Square and square-column transversals are shown. As expected, Eq.(9), increasing square sizes then lead to error decreases among its counterfactual observations at regular and linear rates (lines with different colors). These errors, $\textrm{VAR}_{(k)}$, are depicted by the approximately constant separation between lines. They offer effect estimates under increasing factor variation and EV, Eq.(4). Enumeration of multiple squares, for the same references, reduces errors for individual population members.
F-tests are tests to compare variance of groups and can be used to demonstrate the proposed (ex post) sample stratification strategy. F-tests follow straight-forwardly from the previous error decomposition, Eq.(9) (Appendix.G). Fig.7(m) shows, in particular, results of F-tests for variance of effects for units in cases (a-d) from the same populations (non-stratified data). It shows that F-tests lead to false-negative identification of same effect from variances (outside the statistic $5\%$ significance level) across all cases (horizontal bars are whisker box-plot ranges). The figure also shows the result of tests for subsamples that pass the tests in Eq.(20) (gray), which lead, instead, to true-posititive outcomes in the same simulated data. Fig.7(n) reproduces these results for the non-stationary case, where the correction is more significant. We demonstrate further uses of these estimates in Sect.6.6 Real-World Example. These results are possible due to the limit on effect background association illustrated in Fig.2(b). Fig.7(p-q) shows Bayesian ANOVA regressions, and their confidence intervals (dashed ribbons), for $\textrm{VAR}_{(k)}$ across 1000 runs of unbalanced and correlated cases. Different from the previous, simulations had $3$ fixed levels of unobserved confounders ($q=\{0,10,20\}$), the ground-truth variance in effect observations according to Eq.(9) is shown in black. ANOVA regressions over the full sample (without the proposed combinatorial and gap conditions) are non-convergent. These simple regressions and statistics can help researchers better understand the completeness, and possible generalizability, of their samples and effect estimates.
@@ 6.5 The 80/20 Sample Split
The most universally used rule-of-thumb in ML practice is the 80/20 ratio when splitting samples into train and test subsamples. Although it has remained at the theoretical margins of ML research, largely as application to the problem of overtraining in neural networks, a natural question is: why this proportion? There are two common answers. The first is heuristic, as a consequence of the Pareto Principle. The principle states that ’80% of effects come from 20% of causes’  [Chen et al., 1994]. Causes are used here vaguely, but the principle explains many natural and artificial phenomena. The second answer is more precise, formulated as a scaling-law for the ratio  [Guyon, 1997, Bahri et al., 2021], generalizing  [Amari et al., 1995, Kearns, 1995]. They find that ’the fraction of patterns reserved for the validation set should be inversely proportional to the square root of the number of free adjustable parameters’. In essence, the optimal split is therefore determined by the number of unique factors in a sample, and not its gross number of observations. Eq.(22)(multiple-squares) derived simple bounds, $\tilde{n}$, for sample sizes. The pareto principle and ratio scaling-law offer similar recommendations. Fig.7(o) shows mean difference between recommended sizes, $\tilde{n}$, and the optimal split, $n_{opt}$, across all previous simulation cases. In each case, a sample with $n^{\prime}>100$ observations is divided into 100 random and increasing subsamples. Each subsample adds $n^{\prime}/100$ new observations to the previous and is used as training for the previous supervised methods. The optimal split $n_{opt}$ is taken as the subsample size with an inflection point in the best performance (ACC) across all methods (all samples had ACC peaks, if non-unique the smallest size was taken as optimal). Error is thus $(\tilde{n}-n_{opt})/n^{\prime}$. Although an asymptotic approximation, relating to the asymptotic number of permutations in unbalanced samples, Eq.(22) describes well trade-offs in relation to the EV of training samples. Further empirical illustrations of these combinatorial tradeoffs can be found in  [Ribeiro, 2022a, b]
@@ 6.6 Real-World Example
The COVID19 pandemic continues to threaten the lives and livelihoods of billions around the world. As of the break of the pandemic, there was a rush and expectations for Machine Learning (ML) solutions to help inform policy and individuals’ decisions  [Li et al., 2020, Jin et al., 2021, Verity et al., 2020]. In practice, few ML approaches were truly useful, with SIR semi-deterministic models  [Cramer et al., 2021], or their specializations, often favored. This is, in part, due to the few guarantees current supervised solutions can offer in respect to sample biases and heterogeneity, especially when samples are limited (which coincide with when predictions matter the most). SIR models offer predictions at high aggregate levels - most often country and city levels - taking all individuals therein to be the same. Some sources of heterogeneity have been identified, age being the most obvious. We discuss in particular, how to make predictions with moderate and severe number of unobservabed factors and small $n$. We consider two data sources. The first is the UK Biobank  [Bycroft et al., 2018] (UKB): a dataset with $\sim$500K UK citizens, 100K infected, 5K variables. It includes a wide variety of variables - not only from individuals’ electronic medical records, but also sociodemographic, economic, living, behavior and psychological annotations. The attraction is its completeness, high-dimensionality, and level (individual). We also consider a data dump from the American Center for Disease Control (CDC)  [CDC-Dataset-, 2021] with $\sim$24M cases but limited variables (age, geographic, race, ethnicity, and sex). The interest is as platform to discuss the severe unobserved factors case. Binary outcome is COVID19 infection in both cases. Both datasets include information up to April 1, 2021.
@@
Fig.8(a) shows performance of the previous supervised solutions in the UK Biobank  [Bycroft et al., 2018]. These achieve out-of-sample ACC of $\sim$80with random sampling. We see, under square ordering, a pattern in EV similar to those in the previous tasks. Taking a fourth of the sample leads to 10% increase in ACC over the unseen data. As before, the TSP ordering leads to a sharp decrease in EV in the first-half section. At a fourth of the sample, we have larger confidence over the amount of population predictive coverage the trained models offer. The 10 variables with highest EV, according to Eq.(4), and their biobank codes (parenthesis) were: overcrowded-household (26432), [pop-density (20118), traffic-intensity (24011), time-to-services (26433)], [health-rank (2178), smoking (20116), expiratory-flow (3064), age (24)], job-physical (816), risk-taking (2040). The first is a deprivation index marking individuals in overcrowded households. This is in line with research showing that infection for families happen in a logical-OR fashion (i.e., members multiply their individual exposures to the disease)  [Rader et al., 2020, Emeruwa et al., 2020]. We demonstrated that the presence of many enumerable permutations in samples allow correlated factors, and confounder in particular, to be separated from causes by effect invariance, Principle.1. Variables in brackets have $\rho>0.1$, and are listed in order of effect invariance. There are many squares of size $10$, which indicates that, despite correlations, there are enough observations to also parse out correlated variables’ effects. Population density is favored over neighborhood traffic intensity, and general health over age, for example. Further description of these variables can be found on  [UKB-Showcase-, 2021] from their codes.
@@
Most data used for prediction after the pandemic onset were, however, simpler, with only case counts and a few demographic variables. The COVID19 Case Surveillance database  [CDC-Dataset-, 2021] includes patient-level data reported by US states to the CDC. We use the subsample with simultaneously non-missing age, sex, race, ethnicity and location (county-level). This leads to 10 binary variables and $\sim$11M cases. We additionally generated a set of baseline controls from the corresponding variables in the 2018 American census (county level). Fig.8(c) shows $F$-test estimates for the UKB and CDC samples. Compared to the CDC dataset, the UKB has a comprehensive set of variables and $F$ under $10\%$ significance level. The estimation indicates that the CDC sample has very large numbers of omitted variables. Fig.8(c) also repeats the previous omitted variable plots for the CDC dataset and its 5 counties with highest case counts (top). As expected, $\textrm{VAR}_{(k)}$ ANOVA estimates, Eq.(LABEL:eq-stan), vary significantly across locations (upper-left, with county names) but these have little impact over observed variable pairwise errors (lower lines), as they affect observed square subpopulations uniformly and can be proxied-out, Eq.(9). The perspective suggests that, regardless of the CDC sample’s severe shortcomings, errors $\textrm{VAR}_{(k)}$ can be useful by indicating degree of ’noise’ and EV, Eq.(4) in effect observations. Fig.8(b)(top maps) shows Mean-Squared Error (MSE) of a Leave-One-Out task for the 1489 American counties in the CDC sample, using 3 of the best performers in the previous tasks. Case information from all other counties is used to predict a given location’s COVID19 incidence (addressing the question of which other locations’ infection information could be used to derive optimal estimates for a specific location). MSE is the mean squared difference between the predicted and (held-out) local incidence. Fig.8(b)(lower maps) repeats the task adding $\textrm{VAR}_{(k)}$ to the inputs of these state-of-the-art estimators, which leads to significant gains across locations and algorithms. The lower box-and-whisker plot summarizes results for the Gradient Boosting Machine  [Friedman et al., 2000] (left-most maps). This illustrates that even in clearly confounded samples, where it would be unadvised to assign relative importance to any of its factors, EV estimates can still carry useful information, as error estimates, for subsamples and locations. Other real-world examples are discussed in [Ribeiro, 2022a].

### v5

@@ 7.2 Counterfactual and Correlated Outcome Prediction
While the concepts above can give researchers larger control over the generalizability of learning solutions, we started with the goal of studying the EV of counterfactual predictions. This case is show in Fig.5(b). We take this to be the prediction of effect differences, $\Delta y_{ij}$, from covariate contrasts, $x_{i}-x_{j}$, for all sample unit pairs. The figure illustrates that everyday algorithms are clearly ineffective in this case. The intuition, and why the problem is rarely framed this way, was articulated above: using factor differences leads to samples with increased pairwise overlaps and decreasing permutations, decreasing the capability of algorithms to generalize effectively. The figure shows, however, that models can generalize effectively in samples with the stipulated combinatorial properties, Eq.(14), as in square enumerations. This echoes the increased performance observed over correlated data, and reveals a relationship between these two problems.
@@ 7.3 Causal Effect Estimation and Black-box Explanations
Supervised solutions are highly tuned to generalizability. We now consider implications to causal techniques, which, instead, emphasize biasedness and estimation under non-i.i.d. conditions. Fig.4(d) and Fig.6(d) illustrate two distinct ways to transverse enumerated squares. Each individual square column corresponds to sets of effect observations with a given ordinal rank, and, thus, similar confounding levels and variance, Proposition (1). Distinct square transversals reproduce EV increasing (Definition (2)) and CF non-increasing (Definition (3)) sequences. Imagine we place squares side-by-side, and that each square cell contains a sample unit. We then transverse squares from left-to-right and top-to-bottom. The right-ward sequence of columns is, however, generated in one of two ways: across squares (a square transversal, like in the previous sections) and within squares (a square-column transversal). Using input from these two transversals should make estimators behave asymmetrically in respect to EV and CF. A square transversal corresponds to an order with increasing EV and minimal CF increase. A square-column transversal corresponds to an order with decreasing CF and minimal EV increase. The enumeration production rule for a column $d_{t}$, at time $t$, for each case is also shown in the figures, $d_{1}=1$. Fig.4(c) shows how ACC changes for supervised systems under random, square, and square-column transversals. Fig.6(c) shows how ACC changes, instead, for contemporary causal effect estimators and explaining systems (discussed below). Accuracy is, in this case, sum of squared differences between estimated and ground truth single-variable effects (for all variables). For a set of enumerated squares, for references $x_{0},x_{1},...,x_{t}$ at time $t$, accuracy is calculated for populations $x_{0},x_{1},...,x_{t}$. Lines correspond to mean ACC (and errors) of the best performing algorithm and parameter set (leader) across 1000 runs. Both transversals (square and square-column) supply systems with equally-represented populations but their performance is symmetric in respect to ACC. The random sample ordering strikes a balance between these extremes.
@@
Fig.6(a) shows performance of 3 widely used causal effect estimators and the IME explainer  [Lundberg and Lee, 2017, Lundberg et al., 2019] with random (top) and square-column (bottom) orders for three levels of sample correlation (correlated, square-column transversal)555with $n<100$, omitted datapoints are due to regression non-convergence.. The first estimator is the recent Deconfounder  [Wang and Blei, 2020](Sect. 3.1, linear Bayesian factor model fit with Variational Bayes, logistic outcomes, and Normal priors). The second is g-formula  [Chatton et al., 2020], an efficient solution popular in Epidemiology (Sect. 5.2 Non-parametric Effect Estimation). IME uses the Shapley value calculated from the output of all supervised methods in the last section. A propensity score matching estimator  [Rosenbaum and Rubin, 1983] is included for its popularity and baseline significance in counterfactual solutions. These algorithms (causal and explainer) make different assumptions. The first focuses on biasedness, the second combines (biased and often heuristic) predictive regressions for model selection. As expected, causal effect estimators perform well across correlated cases, compared to the explainer, which makes i.i.d. assumptions, Fig.6(a). Effect estimators loose accuracy with increasing correlation, however, as increasingly less data is uncorrelated in samples and used by these methods.
@@
Squares were used here as means to represent sets of effect observations for a unit $x_{0}$, their permutations in samples, and, finally, to define non-parametric statistics for sample factors’ effects. Our initial definition of effects, Eq.(3), led to consequent definitions and interpretations for the EV, Definition (2), and CF, Definition (3), of counterfactual observations and samples. The unit of analysis here is $\Delta y$, an outcome difference observation, and not the outcomes of regressions. Permutations are not random sample intake orders, but ’observed’ permutations in samples that fulfilled a set of combinatorial and gap conditions, and offer additional guarantees over effect estimates. Eq.(3) thus uses the complete set of $m$-way observed differences in a sample to permute observed counterfactual outcomes and derive effect estimates. The full set of effect background variation, ${\Pi}_{n}(Z\/{a/})$, is the full set of EV-increasing sequences in samples (or their ’external variation’) for individual sample populations.
@@
Two fundamental problems with both the previous explaining and sensitivity analysis approaches are the need for parametric forms for outcomes, and the assumption of i.i.d. variables  [Aas et al., 2021]. The latter is also a problem for standard non-parametric tests and generic permutation-based tests  [Hoeffding, 1948, Yamato and Maesono, 1986]. Fig.6(a) shows that correlations introduce considerable biases into these methods’ estimates of variable importance. Fig.6(b) shows, however, that both methods perform well under square enumerations. Performance increases in both fronts (effect estimation and importance attribution) in this case. In fact, in small samples, the latter dominate recent methods designed specifically for causality and correlated data (upper-right panels). This illustrates that, when samples are limited and effect estimates not externally valid, solutions aimed at generalization (given their inputs are not biased) can also offer gains in causal effect estimation tasks. The Shapley (teal curve) and Sobol indices (black) largely coincide under these conditions, as variance becomes an unbiased indicator of importance and EV, Eq.(4). This illustrates that simple estimators (which do not involve over 20 supervised methods, repeated many times for explanations) can also become effective under square transversals. Notice that both the Deconfounder and g-formula are highly-tuned to DGP assumptions, as both rely on logistic regressions. This exemplifies, in turn, the attraction of non-parametric insights, and that, even when models are correctly specified, there is room for gains with the proposed framework.
@@
Altogether, these results suggest that sample size increase is not the only factor that can lead to increases in the generalizability of algorithms. The number of EV-increasing sequences in the sample play a role. At the same time, the number of CF non-increasing sequences play a role for causal effect estimation (or other tasks related to the estimation of individual factor effect). This observation is in direct contrast with, and suggests significant penalties for, the i.i.d. assumptions of black-box and bootstrap sampling approaches. These methods can observe gains, and control desired statistical tradeoffs, by sampling squares, instead of individual observations.
@@ 7.4 Square Gap and Column Statistics
The panels in Fig.7 (a-l) illustrate the proposed empirical error decomposition for counterfactual effect observations, Eq.(9), in enumerated sample squares. These diagrams depict the mean empirical error among pairs, $(\Delta_{ij}y-\Delta_{uv}y)^{2}$, (y-axis) in their $d$-th column (x-axis). These are observed errors among sample units’ outcomes (in the same square and column), and not effect estimates. The top row, Fig.7(a-d), shows errors in balanced and unbalanced simulations (both transversals). Larger $d\leq m^{\prime}$ lead to pairs exposed to increasing unobserved confounders, and larger variance in effects, for a fixed factor and square reference unit $x_{0}$. Square and square-column transversals are shown. As expected, Eq.(9), increasing square sizes then lead to error decreases among its counterfactual observations at regular and linear rates (lines with different colors). These errors, $\textrm{VAR}_{(k)}$, are depicted by the approximately constant separation between lines. They offer effect estimates under increasing factor variation and EV, Eq.(4).
F-tests are routinely used to compare variance across sample groups and can be used to demonstrate the proposed (ex post) sample stratification strategy. F-tests follow straight-forwardly from the previous error decomposition, Eq.(9) (Appendix.G). Fig.7(m) shows, in particular, results of F-tests for effect variance in units from the same populations in cases (a-d) (non-stratified data). It shows that F-tests lead to false-negative identification of effects with same ground-truth from effect variances (outside the statistic $5\%$ significance level) across all cases (horizontal bars are whisker box-plot ranges). The figure also shows the result of tests for subsamples that pass the tests in Eq.(14) (gray), which lead, instead, to true-posititive outcomes in the same simulated data. Fig.7(n) reproduces these results for the non-stationary case, where the correction is more significant. These results are possible due to the limit on effect background association illustrated in Fig.2(b). Fig.7(p-q) shows Bayesian ANOVA regressions, and their confidence intervals (dashed ribbons), for $\textrm{VAR}_{(k)}$ across 1000 runs of unbalanced and correlated cases. Different from the previous, simulations had $3$ fixed levels of unobserved confounders ($q=\{0,10,20\}$), the ground-truth variance in effect observations according to Eq.(9) is shown in black. ANOVA regressions over the full sample (not passing the proposed combinatorial and gap tests) are non-convergent. This illustrates that, combined with the previous tests, these simple regressions and statistics can help researchers better understand the completeness, and possible generalizability, of their samples and effect estimates.
@@ 7.5 The 80/20 Sample Split
The most universally used rule-of-thumb in ML practice is the 80/20 ratio when splitting samples into train and test subsamples. Although it has remained at the theoretical margins of ML research, largely as application to the problem of overtraining in neural networks, a natural question is: why this proportion? There are two common answers. The first is heuristic, as a consequence of the Pareto Principle. The principle states that ’80% of effects come from 20% of causes’  [Chen et al., 1994]. Causes are used here vaguely, but the principle explains many natural and artificial phenomena. The second answer is more precise, formulated as a scaling-law for the ratio  [Guyon, 1997, Bahri et al., 2021], generalizing  [Amari et al., 1995, Kearns, 1995]. They find that ’the fraction of patterns reserved for the validation set should be inversely proportional to the square root of the number of free adjustable parameters’. In essence, the optimal split is therefore determined by the number of unique factors in a sample, and not its gross number of observations. Eq.(22)(multiple-squares) derived simple bounds, $\tilde{n}$, for sample sizes. The pareto principle and ratio scaling-law offer similar recommendations. Fig.7(o) shows mean difference between recommended sizes, $\tilde{n}$, and the optimal split, $n_{opt}$, across all previous simulation cases. In each case, a sample with $n^{\prime}>100$ observations is divided into 100 random and increasing subsamples. Each subsample adds $n^{\prime}/100$ new observations to the previous and is used as training for the previous supervised methods. The optimal split $n_{opt}$ is taken as the subsample size with an inflection point in the best performance (ACC) across all methods (all samples had ACC peaks, if non-unique the smallest size was taken as optimal). Error is thus $(\tilde{n}-n_{opt})/n^{\prime}$. Although an asymptotic approximation, relating to the asymptotic number of permutations in unbalanced samples, Eq.(22) describes well trade-offs in relation to the EV of training samples.
@@ 7.6 Real-World Example
The COVID19 pandemic continues to threaten the lives and livelihoods of billions around the world. As of the break of the pandemic, there was a rush and expectations for Machine Learning (ML) solutions to help inform policy and individuals’ decisions  [Li et al., 2020, Jin et al., 2021, Verity et al., 2020]. In practice, few ML approaches were truly useful, with SIR semi-deterministic models  [Cramer et al., 2021], or their specializations, often favored. This is, in part, due to the few guarantees current supervised solutions can offer in respect to sample biases and heterogeneity, especially when samples are limited (which coincide with when predictions matter the most). SIR models offer predictions at high aggregate levels - most often country and city levels - taking all individuals therein to be the same. Some sources of heterogeneity were identified, age being the most obvious. We discuss in particular, how to make predictions with moderate and severe number of unobservabed factors and small $n$. We consider two data sources. The first is the UK Biobank  [Bycroft et al., 2018] (UKB): a dataset with $\sim$500K UK citizens, 100K infected, 5K variables. It includes a wide variety of variables - not only from individuals’ electronic medical records, but also sociodemographic, economic, living, behavior and psychological annotations. The attraction is its completeness, high-dimensionality, and level (individual). We also consider a data dump from the American Center for Disease Control (CDC)  [CDC-Dataset-, 2021] with $\sim$24M cases but limited variables (age, geographic, race, ethnicity, and sex). The interest is as platform to discuss the severe unobserved factors case. Binary outcome is COVID19 infection in both cases. Both datasets include information up to April 1, 2021.
@@
Fig.8(a) shows performance of the previous supervised solutions in the UK Biobank  [Bycroft et al., 2018]. These achieve out-of-sample ACC of $\sim$80with random sampling. We see, under square ordering, a pattern in EV similar to those in the simulated data. Taking a fourth of the sample leads to 10% increase in ACC over the unseen data. As before, the TSP ordering leads to a sharp decrease in EV in the first-half. At a fourth of the sample, we have larger confidence over the amount of population predictive coverage the trained models offer. The 10 variables with highest EV, according to Eq.(4), and their biobank codes (parenthesis) were: overcrowded-household (26432), [pop-density (20118), traffic-intensity (24011), time-to-services (26433)], [health-rank (2178), smoking (20116), expiratory-flow (3064), age (24)], job-physical (816), risk-taking (2040). The first is a deprivation index marking individuals in overcrowded households. This is in line with research showing that infection for families happened in a logical-OR fashion (i.e., members multiply their individual exposures to the disease)  [Rader et al., 2020, Emeruwa et al., 2020]. We demonstrated that the presence of many enumerable permutations in samples allow correlated factors, and confounders in particular, to be separated from causes by effect invariance, Principle (1). Variables in brackets have $\rho>0.1$, and are listed in order of effect invariance. There are many squares of size $10$, which indicates that, despite correlations, there are enough observations to also parse out correlated in-sample factors’ effects (that pass out-of-sample gap tests). Population density is favored over neighborhood traffic intensity, and general health over age, for example. Further description of these variables can be found on  [UKB-Showcase-, 2021] from their codes.
@@
Most data used for prediction after the pandemic onset were, however, simpler, with only case counts and a few demographic variables. The COVID19 Case Surveillance database  [CDC-Dataset-, 2021] includes patient-level data reported by US states to the CDC. We use the subsample with simultaneously non-missing age, sex, race, ethnicity and location (county-level). This leads to 10 binary variables and $\sim$11M cases. We additionally generated a set of baseline controls from the corresponding variables in the 2018 American census (county level). Fig.8(c) shows $F$-test estimates for the UKB and CDC samples. Compared to the CDC dataset, the UKB has a comprehensive set of variables and $F$ under $10\%$ significance level. The estimation indicates that the CDC sample has very large numbers of omitted variables. Fig.8(c) also repeats the previous omitted variable plots for the CDC dataset and its 5 counties with highest case counts (top). As expected, $\textrm{VAR}_{(k)}$ ANOVA estimates, Eq.(9), vary significantly across locations (upper-left, with county names) but these have little impact over observed variable pairwise errors (lower lines), as they affect observed square subpopulations uniformly and can be proxied-out, Eq.(9). The perspective suggests that, regardless of the CDC sample’s severe shortcomings, errors $\textrm{VAR}_{(k)}$ can be useful by indicating degree of ’noise’ and EV, Eq.(4) in effect observations. Fig.8(b)(top maps) shows Mean-Squared Error (MSE) of a Leave-One-Out task for the 1489 American counties in the CDC sample, using 3 of the best performers in the previous tasks. Case information from all other counties is used to predict a given location’s COVID19 incidence (addressing the question of which other locations’ infection information could be used to derive optimal estimates for a specific location). MSE is the mean squared difference between the predicted and (held-out) local incidence. Fig.8(b)(lower maps) repeats the task adding $\textrm{VAR}_{(k)}$ to the inputs of these state-of-the-art estimators, which leads to significant gains across locations and algorithms. The lower box-and-whisker plot summarizes results for the Gradient Boosting Machine  [Friedman et al., 2000] (left-most maps). This illustrates that, even in clearly confounded samples, where it would be unadvised to assign relative importance to any factor, EV estimates can still carry useful information, as error estimates, for subsamples and locations. Other real-world examples are discussed in  [Ribeiro, 2022a, b].